OpenStudy (anonymous):

does anyone have a simple formula for working out percentages like 670-35%

OpenStudy (amistre64):

There are generally 3 parts to percentage questions; The Whole, the Part, and the Percent. The Whole = Part/Percent (as a decimal). W=P/D The Part = Percent (as a decimal) * Whole. P=D*P And, Percent (as a decimal) = Part/Whole. D=P/W The Whole amount in this equation is 670. The Part = 670*(.35) 670-(670*.35) = 670-234.5 = 435.5